HUBBLE SPACE TELESCOPE DAILY REPORT #1510 PERIOD COVERED: 0000Z (UTC) 11/02/95 - 0000Z (UTC) 11/03/95 Daily Status Report as of 307/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/1/2 6046 (A Search for Helium Companion of Phi = Persei) The High Resolution Spectrograph (Sides One and Two) was used to = make ultraviolet spectrographic observations of the star Phi Persei (aka, = HD10516) in order to search for a helium companion star. This variable = star is a rapidly rotating Be class star. During the target acquisition = for observation #34, a status buffer error occurred that indicated the = number of fine corrections during x-null balancing exceeded the permitted = maximum. HSTAR 5348 was generated to document this event. Otherwise, the = observations were executed as scheduled, and no further anomalies were = noted. 1.2 Completed WF/PC-2 5977 (Cometary forms in the Helix Nebula) The WF/PC-2 was used to make observations of the Helix Nebula (aka, = NGC-7293, and PK36-57D1E) in several narrow and medium-width filters = looking for comet-like forms. The objective is to obtain sufficient = spectral and spatial information to determine the nature of the comet-like = shapes seen in this planetary nebula. The observations were completed as = scheduled, and no problems were reported. 1.3 Completed WF/PC-2 6106 (The Formation of the Galaxy: Ages for = Globular Clusters in Eridanus) The WF/PC-2 was used to make observations of halo globular clusters = in the constellation of Eridanus. These data will be used to determine = the age of the clusters. The observations were completed as planned, and = no problems were reported. 1.4 Completed WF/PC-2 6187 (WF/PC-2 Cycle 5 Calibration: Earth Flats) The WF/PC-2 was used to make Earth streak flat calibration images. = These streak flats will be used to produce flat fields, to map the OTA = illumination pattern and, in combination with other data, to produce new = superflats. The observations were completed as planned, and no problems = were reported. 1.5 Completed FOS/RD 6167 (FOS Cycle 5: Dark Monitoring) The Faint Object Spectrograph (Red Side) was used to make several = dark calibration images. All exposures were internals. All spectra was = done in the image mode. The observations were completed as planned, and no = problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Operations Notes: There were no significant operational problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-